OpenStudy (precal):

need help with the second derivative of h(t)=2 + cos(2t)

OpenStudy (precal):

h'(t) =-2sin(2t) h"(t)=-4cos(2t)

OpenStudy (anonymous):

Where's your 1st derivative :P

OpenStudy (anonymous):

Good job :)

OpenStudy (precal):

-4cost(2t)=0 divide 4 to both sides cos(2t)=0

OpenStudy (anonymous):

h(t)=2 + cos(2t) h' (t) = - 2* sin 2t h" (t) = -4 cos 2t

OpenStudy (precal):

I was thining of using a double angle formula to solve cos(2t)=cos^2(t)-sin^2(t) then t=pi/4

OpenStudy (precal):



OpenStudy (anonymous):

Is your question about 2nd derivative?

OpenStudy (precal):

I think the second derivative tells me if pi/4 is a max or min, not sure

OpenStudy (precal):

yes, and I was given the interval from - pi to pi

OpenStudy (anonymous):

Max/ min determine by 1st derivative!

OpenStudy (precal):

ok so what does the second derivative tell me??? I am confused. I thought the first derivative also, told me increasing or decreasing. Yes I do recall using that information for max and min

OpenStudy (precal):

ok is it concave up or down or point of inflection?

OpenStudy (anonymous):

Sooryyy, precal, think in reverse side, then you'll get it correctly!

OpenStudy (anonymous):

Yep, 2nd derivative is for inflex, increase/ decreasing!

OpenStudy (precal):

then when I solve it, I think it is pi/4 what happens at that point according to the 2nd derivative

OpenStudy (precal):

well I still not sure what to do

OpenStudy (anonymous):

That's the whole point? I still have no clue what your question about? What your purpose of taking second derivative?

OpenStudy (anonymous):

If you find min/ max -> let f'(x) = 0

OpenStudy (precal):

ok I think the question is: Determine the intervals on which the following functions are concave up or concave down. Identify any inflection points. I am helping someone understand the problem. I think they mislead me to do the 2nd derivative.

OpenStudy (anonymous):

The purpose of 2nd derivative is find inflex, increase => that what they want

OpenStudy (anonymous):

If so let f"(x) = 0